A chemist must prepare 350ml of a chemical solution ,itmis to be made up of 2 parts of alcohol and 3 parts of acid .how much of each should be used?

Grade:6

1 Answers

Arun
5 years ago

Let n be the number of milliliters in each part. The picture above shows the solution. From the diagram, we have

2n+3n = 350

5n = 350

n = 350/5 = 70

But n = 70 is not the answer to the original problem. Each part has 70 ml.

The amount of alcohol is 2n = 2(70) = 140, and the amount of acid is 3n = 3(70) = 210.

Thus, the chemist should use 140ml of alcohol and 210ml of acid.